If a 65 mL solution of HCl has a concentration of 0.326 M, what is its pH? Also, how many
grams of HCl are in the solution?
1
Expert's answer
2020-05-03T15:12:44-0400

Solution:

pH = - lg[HCl] = 0.487

m(HCl) = n(HCl)*M(HCl) = C(HCl)*V(HCl)*M(HCl) = 0.326 M*0.065L*36.46g/mol = 0.77 g